There are 200 data of 1 and 800 data of 0.

Sample proportion is 0.4

Assume the population proportion is 0.2. You take a sample of 200.

1. Calculate the obtained z statistic. Find it on the output.

2. What is the probability of getting the sample proportion or less? Find it on the output.

3. Sketch the sampling distribution of the proportion that assumes the alternative hypothesis is true. Label: The expected proportion, observed proportion of your sample and the standard error off the sampling distribution.
Answer #1

1.

Test Statistic under Ho: Pe 0.4-0.2 0.20.8 = --=7.071068 * 200

2. P(P<0.4) = P(Z < 7.071068) = 0.0000

3. The expected proportion is 0.2

The SE of the smapling distribution is SQRT(0.2*0.8/200) = 0.028284

Here P-value = 0.000 < alpha .05 so we reject H0

Thus we conclude that the population proportion is not equal to 0.2